Calculate the required rate of return for Food Inc. Assuming that (1) investors expect a 2.0% rate of inflation in the future, (2) the real risk-free rate is 3.5%, (3) the market portfolio return is 7.5%, (4) the firm has a beta of 2.00, and (5) its realized rate of return has averaged 12.0% over the last 5 years. (Hint: You will need to get the market premium first in the CAPM model).
Answer #1

Market premium = beta *(rf-rm)

= 2*(7.5-3.5)

= 8

CaPM model = rf+ market premium

= 3.5+8 = 11.5% is the market return

There is an inflation of 2%

So, required rate of return is 11.5%+2% = 13.5%
Answer #2

Market premium = beta *(rf-rm)

= 2*(7.5-3.5)

= 8

CaPM model = rf+ market premium

= 3.5+8 = 11.5% is the market return

There is an inflation of 2%

So, required rate of return is 11.5%+2% = 13.5%